# GRPO-MA: Multi-Answer Generation in GRPO for Stable and Efficient Chain-of-Thought Training

### v2

@@ 4.2.2 Variance of the Thought Advantage
$\mathrm{Var}[A(th_{i})]\;\approx\;\frac{1}{M\,\sigma_{\mu_{R}}^{2}}\sum_{k=1}^{K}\Biggl(\;\delta_{ik}-\tfrac{1}{K}-\tfrac{\tilde{\mu}_{i}\,\tilde{\mu}_{k}}{K-1}\;\Biggr)^{2}\sigma_{R_{k}}^{2}$ (4)
@@
For the full derivation on the variance of though advantages and answer advantages, please refer to Appendix A.2.
@@ 4.2.3 Analysis of the Variance Structure
This expression clearly shows that the variance of the thought advantage is inversely proportional to $M$. Increasing the number of answers sampled per thought directly reduces the variance of the thought’s value estimate, leading to more stable thought’s advantage estimation.
Additionally, let us analyze how variance changes as $K$ increases. As $K\to\infty$, by the law of large numbers, the sample variance of the true thought values $\sigma_{\mu_{R}}^{2}=\tfrac{1}{K-1}\sum_{k=1}^{K}(\mu_{R_{k}}-\mu_{\bar{R}})^{2}$ converges to the population variance $\sigma_{\pi}^{2}=\mathrm{Var}[\mu_{R_{i}}]$, which characterizes the variability of the true thought values $\mu_{R_{i}}$ across the population of thoughts sampled from the distribution $\pi_{\theta}(\cdot\mid p)$. Next, we analyze the sum by splitting it based on the Kronecker delta, $\delta_{ik}$. The single term where $k=i$ converges to a non-zero constant, since expressions like $(\delta_{ik}-\tfrac{1}{K}-\dots)$ approach $1$. Conversely, the sum of the other K1K-1 terms, where $k\neq i$, vanishes because each term is of order $O(1/K^{2})$, making their total sum $O(1/K)$. This leads to the final result where the variance converges to a limit determined only by the properties of the thought $th_{i}$ itself: $\lim_{K\to\infty}\mathrm{Var}[A(th_{i})]=\tfrac{\sigma_{R_{i}}^{2}}{M\,\sigma_{\pi}^{2}}$.
The analysis of increasing $K$ and $M$ indicates that while increasing $K$ alone preserves a minimum variance , increasing $M$ allows the variance to decrease continuously and approach 0.
@@ 5 Experiments
We evaluate GRPO-MA on Math Yu et al. (2025), Code PrimeIntellect (2024); White et al. (2024), several distinct vision tasks (Object Detection contributors (2024), Affordance Prediction Myers et al. (2015); Luo et al. (2022), Trajectory Prediction Ji et al. (2025), Demand Prediction Wang et al. (2024), OCR-based VQA Biten et al. (2019); Tito et al. (2021)) and a Simulator-based Visual Manipulation task Li et al. (2024). Our experiments use Qwen2.5-VL-3B-Instruct Bai et al. (2025) as the base model, with all training conducted on four H100 80G GPUs using LoRA Hu et al. (2022) for parameter-efficient fine-tuning. For each task, we conduct a group of experiments separately.
We introduce the TKKAMM notation to unify the representation of the GRPO and GRPO-MA methods. In this notation, $K$ represents the number of thoughts, and $M$ denotes the number of answers generated per thought. The notation corresponds to GRPO when $M=1$ and to GRPO-MA when $M>1$. For instance, T4A4 signifies a process of generating 4 thoughts, with each thought producing 4 answers, resulting in a total of 16 responses.
More details (datasets, hyperparameters, training settings) are in the appendix A.3 and  A.4.
@@ 5.1.1 Task Setting and Metric
Math Given a math problem, output the correct solution. Metric: pass@10/32 Chen et al. (2021).
Code Given a programming problem, output the solution code. Metric: pass@10/32.
Object Detection Given an image and a specified object name, output bounding boxes. Metric: proportion of predictions with Intersection-over-Union (IoU) above threshold.
Affordance Prediction Given an image and an affordance (e.g., grasping, holding), output 2D coordinates. Metric: proportion of matched points.
Trajectory Prediction Given an image and a manipulation instruction, output the 2D end-effector trajectory. Metrics: Discrete Fréchet Distance (DFD) Eiter et al. (1994), Hausdorff Distance (HD) Huttenlocher et al. (2002), Root Mean Square Error (RMSE), EndPoint Distance.
Demand Prediction Given an image and a human demand instruction, output 2D coordinates of the demanded object. Metric: proportion of correct points.
OCR-based VQA Given an image and a question requiring text understanding (e.g., infographics, scene text, documents), output the answer. Metric: Average Normalized Levenshtein Similarity (ANLS) Biten et al. (2019).
Please note that, unlike the <think> and <answer> tags used in GRPO, we design a distinct structured output format for Math consisting of three tags: <analysis>, <process>, and <answer>. Multi-sampling is applied to both <process> and <answer>. For other tasks, Multi-sampling is applied only to the <answer>.
We track the Gradient Spike Score (GSS) Huang et al. (2025a) to measure gradient stability, defined as $\text{GSS}(g_{i})=\tfrac{|g_{i}|}{\tfrac{1}{T+1}\sum_{j=0}^{T}|g_{j}|}$, where $g_{j}$ represents the gradient at the j-th time step. We report the number of spikes above 10 (GSS@10), where smaller is better. For all tasks, we also report the per-step training time (s).
@@ 5.1.2 Baselines
We adopt models from the Qwen2.5-VL-Instruct series (3B, 7B, and 72B) Bai et al. (2025) as baselines to evaluate the performance of general-purpose models on our tasks. In addition, we train Qwen2.5-VL-3B-Instruct with real labels using supervised fine-tuning (SFT) to compare against GRPO, denoted as SFT in the results. Finally, we compare our proposed GRPO-MA with GRPO under different numbers of responses to demonstrate the superiority of GRPO-MA in terms of training efficiency and performance.
5.1.3 Main Results
The experimental results are presented in Tab. 1 (Math and Code Problem), Tab. 2 (Object Detection, Affordance Prediction and Demand Prediction), and Tab. 3 (OCR-based VQA and Trajectory Prediction). Across multiple visual tasks, our proposed GRPO-MA outperforms both the GRPO and SFT under various settings, demonstrating its excellent versatility across diverse tasks. Compared to T4A1, T4A4 achieves significant performance gains with only about a 15% increase in training time. Compared to T16A1, T4A4 achieves comparable or even slightly better performance with about a 40% reduction in training time, which demonstrates that GRPO-MA does not involve a trade-off between training efficiency and training performance, but rather enhances both simultaneously.
@@ 5.1.3 Main Results
Model TN AN S/S GSS Pass@10 Pass@32 S/S GSS Pass@10 Pass@32
@@
Gradient Stability In most experiments, T4A4 achieves the lowest GSS@10, indicating the best gradient stability during training, consistent with our theoretical analysis: as a crucial component of gradient magnitude, the more stable estimation of the advantage value also contributes to greater gradient stability.
Case Study As illustrated in Fig. 2, we present a case study to contrast the reasoning processes of T4A4 and T4A1 for the object detection task. T4A4 focuses on the general vicinity of the target object and its surrounding context. Conversely, T4A1 fails to detect the target, instead paying its attention on the central region of the image. Additional case studies are provided in the appendix A.5.
@@ 5.2.1 Task Setting
We adapt most of the experimental settings introduced in ManipLLM Li et al. (2024), which provides a simulator-based framework for visual manipulation tasks. To increase the difficulty of the task, we introduce two modifications to the experimental setup. First, to ensure greater observational diversity, the camera is reconfigured to view the target object from a randomly sampled angle in each trial. Second, we adapt a stricter success criterion: an attempt is immediately deemed a failure if the predicted contact point does not lie on the surface of the target object. Following ManipLLM, when the model outputs a grasping point on the image, we execute a rule-based grasping strategy. Specifically, the sucker approaches along the surface normal at the predicted point, and the subsequent trajectory is adjusted depending on the object category. For evaluation, we report the proportion of predicted points that lead to successful manipulation. Through data collection and training, we observe that this task is highly reward-sparse, since solving it requires the model to reason about object-specific interaction dynamics.
5.2.2 Baselines
We adapt some of the same baselines used in visual tasks and added several additional baselines: ManipLLM-7B, CoT-SFT, and GRPO-NoThink.
ManipLLM-7B They collects a large number of successful samples in the simulator and constructs multiple task-specific question-answer pairs, utilizing the SFT training approach. We have fine-tuned their weights in the new settings.
CoT-SFT We collect successful samples of GRPO-MA-T4A4 (including the chain of thoughts and answers), then fine-tune Qwen2.5-VL-3B using SFT.
GRPO-NoThink We employ GRPO to train the Qwen2.5-VL-3B, but we do not require the model to generate a thought process; instead, it directly produces the answers.
5.2.3 Main Results
@@
The experimental results are presented in Tab. 4. A direct comparison reveals that the performance of T4A4 is significantly superior to that of T4A1. This outcome demonstrates that in tasks with extremely sparse rewards, such as multi-modal manipulation, employing a multi-answer sampling strategy leads to a more stable training process and facilitate sampling of effective responses.
Furthermore, our experiments provide valuable insights into the indispensable role of the Chain of Thought (CoT) in this context. We observe that the GRPO-NoThink model, which ablates the CoT while sampling an equal number of answers as GRPO-MA-T4A4, suffers a substantial degradation in performance. This result, along with the strong performance of the CoT-SFT model, clearly indicates that a high-quality CoT is a critical prerequisite for generating superior answers and effectively tackling such complex tasks.
@@ 5.3 Ablation Study
We conduct a detailed ablation study on the trajectory prediction task to analyze the effect of the number of generated answers $M$ per thought, as shown in Fig. 3. The results indicate that as $M$ increases, all evaluation metrics decrease, although the rate of decline becomes progressively smaller.
Surprisingly, T4A3 features 4 thoughts and 12 answers, outperforming T16A1’s 16 thoughts and 16 answers across all metrics. One possible explanation for this finding is that the importance of reward signal richness (the number of answers) is less significant than the quality of thoughts; filtering out higher-quality thoughts has a greater impact on the overall training process. Specifically, our method assesses a thought’s quality by averaging the rewards of its $M$ subsequent answers ($V(th_{i})=\tfrac{1}{M}\sum_{j=1}^{M}R_{i,j}$). With $M=3$, T4A3 obtains a more stable and reliable estimate of each thought’s value, effectively reducing the noise from any single-answer evaluation. In contrast, T16A1’s approach ($M=1$) is far more susceptible to randomness, as a single, potentially noisy reward is used to judge the entire thought.
5.4 Inconsistency Analysis
We quantify the inconsistency between thoughts and answers during training. For a thought $th_{i}$ with $M$ answers, if $\operatorname{sign}(A(th_{i}))\neq\operatorname{sign}(A(ans_{i,j}))$, we mark it as inconsistent. The inconsistency rate is defined as $\text{InconsistencyRate}=\tfrac{1}{KM}\sum_{i=1}^{K}\sum_{j=1}^{M}1[A(th_{i})A(ans_{i,j})<0]$, where $\bm{1}[\cdot]$ denotes the indicator function, which equals $1$ if the condition inside holds and 0 otherwise.
Accuracy reward curves and richness of reward signal analysis are in the appendix A.6.

### v3

@@ 4.2.2 Approximate Variance Estimation
$\boxed{\mathrm{Var}[A(th_{i})]\;\approx\;\frac{1}{M\,\sigma_{\mu_{R}}^{2}}\sum_{k=1}^{K}\Biggl(\;\delta_{ik}-\tfrac{1}{K}-\tfrac{\tilde{\mu}_{i}\,\tilde{\mu}_{k}}{K-1}\;\Biggr)^{2}\sigma_{R_{k}}^{2}}$ (3)
@@
For the full derivation on the variance of though advantages and answer advantages, please refer to Appendix C.
4.2.3 Controllable and uncontrollable sources of variance.
Apart from the sampling dimensions $K$ and $M$, the remaining terms in Eq. (3) depend on the particular sampled thoughts and vary stochastically across iterations. Within the GRPO-style estimator considered here, these quantities are not directly controllable by the learning algorithm. To the best of our knowledge, there is currently no stable and generally applicable method to control them in GRPO-style training. Even if such methods were to exist, they would be complementary to answer-level branching rather than a substitute. Accordingly, $K$ and $M$ constitute the primary controllable parameters for variance control in our analysis.
@@ 4.2.4 Analysis of the Variance Structure
Combining the above results, we now compare the effects of increasing the number of sampled thoughts $K$ versus increasing the number of answers $M$ for each thought.
As $K\to\infty$, by the law of large numbers, the sample variance of the true thought values,
$\sigma_{\mu_{R}}^{2}=\frac{1}{K-1}\sum_{k=1}^{K}(\mu_{R_{k}}-\mu_{\bar{R}})^{2},$
converges almost surely to the population variance
$\sigma_{\pi}^{2}=\mathrm{Var}_{\,th\sim\pi_{\theta}(\cdot\mid p)}\!\bigl[E[V(th)\mid p,th]\bigr].$
which characterizes the intrinsic variability of true thought values across the population induced by the policy $\pi_{\theta}$ and prompt $p$.
To analyze the remaining terms in the variance expression, we decompose the relevant summation according to the Kronecker delta $\delta_{ik}$. The single term with $k=i$ converges to a non-zero constant, since expressions of the form $(\delta_{ik}-\tfrac{1}{K}-\cdots)$ approach $1$ as $K\to\infty$. In contrast, each of the remaining K1K-1 terms with $k\neq i$ is of order $O(1/K^{2})$, and therefore their total contribution vanishes at rate $O(1/K)$.
As a result, for any fixed number of answers $M<\infty$, the variance of the advantage estimator converges to a strictly positive limit:
$\lim_{K\to\infty}\mathrm{Var}[A(th_{i})]=\frac{\sigma_{R_{i}}^{2}}{M\,\sigma_{\pi}^{2}}.$
This limit depends only on the intrinsic response variance $\sigma_{R_{i}}^{2}$ of the thought $th_{i}$ and the population-level variance $\sigma_{\pi}^{2}$, and cannot be further reduced by increasing $K$ alone when $M$ is held fixed.
In contrast, for any fixed number of sampled thoughts $K<\infty$, increasing the number of answers $M$ directly suppresses the variance of the estimated thought advantage. In this regime, the variance scales inversely with $M$, and thus
$\lim_{M\to\infty}\mathrm{Var}[A(th_{i})]=0,\ \mathrm{with}\ \mathrm{Var}[A(th_{i})]=O\!\left(\frac{1}{M}\right).$
These results reveal a fundamental asymmetry in thought-level advantage estimation under GRPO-style training without value models. Increasing the number of sampled thoughts $K$ yields no predictable variance reduction at finite $K$ and is only guaranteed to converge to a strictly positive variance floor as $K\to\infty$. In contrast, increasing the number of answers per thought $M$ induces a monotonic and controllable reduction in variance, decreasing it to zero as $M$ increases. As a result, effective variance reduction cannot be achieved through thought expansion alone. Absent additional mechanisms for directly controlling thought content, tree-style branching that induces multiple answers per thought becomes potentially necessary for stabilizing thought advantage estimation.
@@ 5 Experiments
We evaluate GRPO-MA on Math (Yu et al., 2025), Code (PrimeIntellect, 2024; White et al., 2024), several distinct vision tasks (Object Detection (contributors, 2024), Affordance Prediction (Myers et al., 2015; Luo et al., 2022), Trajectory Prediction (Ji et al., 2025), Demand Prediction (Wang et al., 2024), OCR-based VQA (Biten et al., 2019; Tito et al., 2021)) and a Simulator-based Visual Manipulation task (Li et al., 2024). Our experiments use Qwen2.5-VL-3B-Instruct (Bai et al., 2025) as the base model, with all training conducted on four H100 80G GPUs using LoRA (Hu et al., 2022) for parameter-efficient fine-tuning. For each task, we conduct a group of experiments separately.
More details (tasks, datasets, hyperparameters, training settings) are in the appendix D and  E.
@@ 5.1.1 Task Setting and Metric
Table 1 summarizes evaluation metrics for text and vision tasks. These tasks collectively cover a wide range of modalities and reasoning types, enabling a comprehensive evaluation of GRPO-MA.
Table 1: Evaluation metrics for each task.
Task Evaluation Metric
Math Pass@k  (Chen et al., 2021b)
Code Pass@k
Object Detection Accuracy: proportion of predictions with IoU >> threshold.
Affordance Prediction Accuracy: proportion of correctly matched points.
Trajectory Prediction DFD (Eiter et al., 1994), HD (Huttenlocher et al., 2002), RMSE, EndPoint Dist.
Demand Prediction Accuracy: proportion of correct points.
OCR-based VQA ANLS (Biten et al., 2019)
For Math, we adopt a structured output format using <analysis>, <process>, and <answer> tags. GRPO-MA applies multi-sampling on both <process> and <answer>. For all other tasks, multi-sampling is applied only to <answer>.
We track the Gradient Spike Score (GSS) (Huang et al., 2025a) to measure gradient stability, defined as $\text{GSS}(g_{i})=\tfrac{|g_{i}|}{\tfrac{1}{T+1}\sum_{j=0}^{T}|g_{j}|}$, where $g_{j}$ represents the gradient at the j-th time step. We report the number of spikes above 10 (GSS@10), where smaller is better.
@@ 5.1.2 Baselines
We adopt models from the Qwen2.5-VL-Instruct series (3B, 7B, and 72B) (Bai et al., 2025) as baselines to evaluate the performance of general-purpose models on our tasks. In addition, we train Qwen2.5-VL-3B-Instruct with real labels using supervised fine-tuning (SFT) to compare against GRPO, denoted as SFT in the results. Finally, we compare our proposed GRPO-MA with GRPO under different numbers of responses to demonstrate the superiority of GRPO-MA in terms of training efficiency and performance.
5.1.3 Main Results on Text and Vision Task
@@
Model TN AN S/S GSS Pass@10(N=100) Pass@32(N=100) S/S GSS Pass@10(N=100) Pass@32(N=100)
@@
The experimental results are presented in Tab. 2 (Math and Code), Tab. 3 (Object Detection, Affordance Prediction, and Demand Prediction), and Tab. 4 (OCR-based VQA and Trajectory Prediction). Across multiple tasks, our proposed GRPO-MA consistently outperforms both GRPO and SFT under various settings, demonstrating strong versatility across diverse domains. Please see Appendix G.6 for additional Math and Code results on GSM8K (Cobbe et al., 2021) and HumanEval (Chen et al., 2021a). Case studies are included in Appendix F.
Gradient Stability. While our theoretical analysis focuses on the variance of thought-level advantage estimation, its implications for optimization stability follow from the structure of the GRPO-MA objective in Eq. 2, where advantage estimates play a central role in shaping gradient updates. By reducing advantage variance through answer-level branching, GRPO-MA indirectly suppresses extreme gradient magnitudes and leads to more stable optimization. Consistent with this implication, T4A4 achieves the lowest GSS@10 in most experimental settings, indicating a substantially reduced frequency of gradient spikes during training. Additional grad_norm and GSS curves are provided in Appendix G.5.
[…]
These results indicate that the efficiency gains of GRPO-MA do not arise from trading off performance for train speed. Instead, they stem from a more effective allocation of additional samples across sampling levels: by concentrating extra sampling on the answer dimension, GRPO-MA yields more stable advantage estimates, allowing improved optimization behavior to translate into both faster training and stronger performance. Crucially, this also implies that such gains cannot be achieved by expanding thought samples alone, rendering answer-level branching a necessary design choice for achieving reliable variance reduction and efficient optimization under GRPO-style training.
@@ 5.2.1 Task Setting
We largely follow the experimental protocol of ManipLLM (Li et al., 2024), a simulator-based framework for visual manipulation. To make the task more challenging, we introduce two changes. First, to increase observational diversity, the camera viewpoint is randomly sampled in each trial. Second, we adopt a stricter success criterion: an attempt is immediately counted as a failure if the predicted contact point is not on the target object surface. As in ManipLLM, when the model predicts a grasp point in the image, we execute a rule-based grasping routine. The suction cup approaches the predicted point along the local surface normal, and we further adjust the subsequent trajectory based on the object category. We evaluate performance by the success rate of predicted points that result in successful manipulation.
@@
5.2.2 Baselines
We adapt some of the baselines used above and add several additional baselines. ManipLLM-7B: They collects a large number of successful samples in the simulator and constructs multiple task-specific question-answer pairs, utilizing the SFT training approach. We fine-tune their weights in the new settings. CoT-SFT: We collect successful samples of GRPO-MA-T4A4 (including the chain of thoughts and answers), then fine-tune Qwen2.5-VL-3B using SFT. GRPO-NoThink: We employ GRPO to train the Qwen2.5-VL-3B, but we do not require the model to generate a thought process; instead, it directly produces the answers.
5.2.3 Main Results on Manipulation Task
Manipulation tasks with sparse rewards constitute an extreme high-variance regime for GRPO-style training [^2] For simplicity, we assume a binary reward $R\in\{0,1\}$ that follows a Bernoulli distribution with success probability $p$. When $p$ is very small, reward observations are extremely noisy, making single-sample estimates unreliable and leading to high variance in thought-level advantage estimation., where instability in thought-level advantage estimation is particularly pronounced. The experimental results are presented in Tab. 5. A direct comparison shows that T4A4 significantly outperforms T4A1, indicating that answer-level branching is especially effective in this setting by stabilizing training and facilitating the discovery of successful responses under sparse supervision. This result further supports our variance-based analysis, suggesting that tree-style branching becomes particularly important for reliable credit assignment in high-variance manipulation tasks.
5.3 More Baseline Comparison
Table 6: More Baseline Comparison on Trajectory Prediction .
Model DFD HD RMSE EP
GRPO 187.99 172.58 140.80 142.74
GRPO-CARE 188.23 170.82 139.83 144.85
DAPO 184.08 167.75 136.18 146.99
Dr.GRPO 180.56 165.95 135.86 140.05
TreeRPO-128 165.27 149.77 123.42 131.45
TreeRPO-64 157.03 144.52 113.99 125.08
GRPO-MA 151.10 138.29 111.59 120.60
To provide a broader comparison, we evaluate four recent GRPO-based variants—GRPO-CARE, DAPO, Dr.GRPO, and TreeRPO—on the trajectory prediction task. As shown in Table 6, GRPO-MA achieves the best performance across all metrics. GRPO-CARE, DAPO, and Dr.GRPO improve GRPO via loss or normalization refinements but do not introduce explicit branching, so thought-level advantages are still estimated from single outcomes and remain high-variance under sparse rewards. TreeRPO relies on token-length–based branching, where the suffix (e.g., -64 or -128) denotes the maximum segment length used for splitting responses. As a result, its behavior is sensitive to response length: TreeRPO-128 degenerates to standard GRPO when responses are short, resulting in poorer performance compared to TreeRPO-64. Although TreeRPO-64 mitigates this issue and improves performance relative to TreeRPO-128, it still underperforms GRPO-MA, which we conjecture is due to higher-variance estimation of values (or advantages) for thoughts in sub-branches induced by branching at intermediate segments. In contrast, our XML-tag–based branching operates on semantic boundaries, enabling stable branching and reducing variance in thought-level advantage estimation across variable-length outputs.
5.4 The Effectiveness on Larger Model Size
To study how GRPO-MA behaves as model capacity increases, we conduct a scaling experiment on Qwen2.5-VL-7B-Instruct. The 7B scale is a commonly used “large” checkpoint for tree-search- or multi-sampling-based reasoning methods (Guo et al., 2025; Kazemnejad et al., 2024; Li et al., 2025), offering a practical trade-off between model capacity and the computation required by expanded rollouts. We compare GRPO-MA with a matched GRPO baseline using identical training settings on the trajectory prediction task. Results are summarized in Table 7. Notably, GRPO-MA still outperforms GRPO, indicating that multi-answer sampling continue to benefit larger models.
Table 7: Qwen2.5-VL-7B-Instruct on Trajectory Prediction.
Model DFD HD RMSE EP
7B-T8A1 167.32 152.34 129.50 134.60
3B-MA-T4A4 151.10 138.29 111.59 120.60
7B-MA-T8A4 134.67 121.89 103.64 103.27
5.5 The Effectiveness on Different Model Architecture
we evaluate GRPO-MA on the pure-text language model rather than a vision language model, Qwen3-4B-Instruct, for math and code reasoning tasks. The training settings remain the same as before. These results shown in Table 8 indicate that GRPO-MA continues to provide improvements even on pure text models, Qwen3-4B-Instruct, supporting the generality of our approach on different model architectures.
Table 8: Qwen3-4B-Instruct on Math and Code.
Model Pass@1(N=10) on Math Pass@1(N=10) on Code
GRPO-T4A1 17.53 4.22
GRPO-MA-T4A4 18.07 6.25
@@ 5.6 Ablation Study
We conduct a detailed ablation study on the trajectory prediction task to analyze the effect of the number of answers per thought $M$, as shown in Fig. 2. The results indicate that as $M$ increases, most evaluation metrics tend to decline, although the rate of decline becomes progressively smaller.
Surprisingly, T4A3 (4 thoughts, 12 answers) outperforms T16A1 (16 thoughts, 16 answers) across all metrics. A plausible reason is that thought quality matters more than reward-signal quantity: we score each thought by averaging the rewards of its $M$ sampled answers, $V(th_{i})=\tfrac{1}{M}\sum_{j=1}^{M}R_{i,j}$. With $M=3$, T4A3 yields a more stable value estimate than T16A1’s single-sample evaluation ($M=1$), which is more sensitive to reward noise.